An airplane flew to San Diego with a tailwind and averaged 485 miles per hour. On the return trip the plane only averaged 275 miles per hour. find the speed of the wind and the speed of the plane in still air.
Answer by mananth(16946)

x-- speed of plane
y--- speed of wind
with tail wind
speed = x+y= 485
against wind
x-y= 275
solve for x &y
add the equations
2x= 760
/2
x= 380 ----speed of plane
y=485-380
y=105 mph